Mary is 9 years younger than Anne. Tom is twice as old as Anne. If the sum of their ages is 95, how old are Mary, Tom, and Anne?
Answer by ikleyn(52788)

.
Let x be the Anne's age.


Then the Mary's age is (x-9) years and Tom's age is 2x years.


The total is


    x + (x-9) + 2x = 95.


Simplify and solve for x.


    4x - 9 = 95,

    4x     = 95 + 9

    4x     = 104

     x     = 104/4 = 26.


ANSWER.  Anne is 26 years old;  Mary is 26-9 = 17 years old;  Tom is  26*2 = 52 years old.


CHECK.   26 + 17 + 52 = 95 years.    ! Correct !